OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 2x - 3y = -18. x intercept = (9, 0); y intercept = (0, -6) x intercept = (6, 0); y intercept = (0, -9) x intercept = (-9, 0); y intercept = (0, 6) x intercept = (-6, 0); y intercept = (0, 9)

OpenStudy (anonymous):

I think its D

jimthompson5910 (jim_thompson5910):

to find the x-intercept, plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

2x - 3y = -18 2x - 3(0) = -18 2x - 0 = -18 x = ???

OpenStudy (anonymous):

i am sorry i sssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssm